OpenStudy (anonymous):

How to do negative powers? Like if i had 6 to the power of -4

OpenStudy (anonymous):

6^(-4) = 1/6^(4) Power Law states that 2^(-1) = 1/(2^(1))

OpenStudy (mertsj):

That negative exponent means "the reciprocal of" so 6^-4 means the reciprocal of 6 to the 4th power.

OpenStudy (anonymous):

So the answer would be 4/6? and what is a reciprocal?

OpenStudy (mertsj):